# Bayesian Unsupervised Disentanglement of Anatomy and Geometry for Deep Groupwise Image Registration

### v2

@@ 1 Introduction
[…]
@@
$L$ the number of levels of latent variables $N$ the number of images in the group
@@ 2.1 Groupwise Image Registration
[…]
@@ 2.2 Deep Feature-Based Image Registration
[…]
@@ 3.1 Hierarchical Bayesian Inference
$\mathcal{L}(\bm{\theta},\bm{\psi}\mid\bm{u})$ $&\mathbb{E}_{q(\bm{z},\bm{v}\mid\bm{u};\bm{\psi})}[\log p(\bm{u}\mid\bm{z},\bm{v};\bm{\theta})]\\ &-D_{\text{KL}}\big{[}q(\bm{z},\bm{v}\mid\bm{u};\bm{\psi})\;\|\;p(\bm{z})p(\bm{v})\big{]},\end{aligned}$ (6)
@@
[…]
@@ 3.2.1 Structural Representations for Variable Group Sizes
$=$ $\frac{1}{2}\Bigg{[}\log\frac{\absolutevalue{\bm{\Sigma}_{\bm{z},j}^{\diamond}}}{\absolutevalue{\bm{\Sigma}_{\bm{z}}^{*}}}+\operatorname{tr}\left(\bm{\Sigma}_{\bm{z},j}^{\diamond\,-1}\bm{\Sigma}_{\bm{z}}^{*}\right)$
$\quad\;+(\bm{\mu}_{\bm{z},j}^{\diamond}-\bm{\mu}_{\bm{z}}^{*})^{\intercal}\bm{\Sigma}_{\bm{z},j}^{\diamond\,-1}(\bm{\mu}_{\bm{z},j}^{\diamond}-\bm{\mu}_{\bm{z}}^{*})\Bigg{]}+\text{const.}$.
@@
Here, we extend the framework using categorical distribution to improve the interpretability of the structural representations. Particularly, we propose to model them by independent categorical latent variables, i.e., $q_{j}^{\diamond}(\bm{z}\mid u_{j},\bm{v}_{j};{\bm{\psi}_{j}})=\operatorname{Cat}(\bm{z};\bm{\pi}_{j,1}^{\diamond},\dots,\bm{\pi}_{j,K}^{\diamond})$, where $\bm{\pi}_{j,k}^{\diamond}=(\pi_{\bm{\omega},j,k}^{\diamond})_{\bm{\omega}\in\Omega}\in[0,1]^{\absolutevalue{\Omega}}$ and $\sum_{k=1}^{K}\bm{\pi}_{j,k}^{\diamond}=\bm{1}\in\mathds{R}^{\absolutevalue{\Omega}}$. Thus, the geometric mean becomes another categorical distribution
@@
$\bm{\pi}_{k}^{*}=\frac{\left[\prod_{j=1}^{N}\bm{\pi}_{j,k}^{\diamond}\right]^{\nicefrac{{1}}{{N}}}}{\sum_{k=1}^{K}\left[\prod_{j=1}^{N}\bm{\pi}_{j,k}^{\diamond}\right]^{\nicefrac{{1}}{{N}}}}\in[0,1]^{\absolutevalue{\Omega}},$
@@ 3.2.2 Categorical Reparameterization Using Gumbel-Rao
Unfortunately, the Gumbel-Max trick [75] for sampling discrete random variables is not differentiable w.r.t. its parameterization, making it unsuitable for stochastic gradient estimators. Concurrent works [76, 77] have then introduced the Gumbel-Softmax (GS) distribution as a continuous relaxation of the discrete categorical variable, which admits a biased reparameterization gradient estimator. Here, we use a variant of the GS estimator, namely the Gumbel-Rao (GR) estimator [78], to further reduce the variance in gradient estimation via Rao-Blackwellization.
@@ 4.2 Inference of Velocity Fields
$\sigma_{\bm{\varphi}_{j}^{l}}^{2}=\frac{1}{\absolutevalue{\Omega^{l}}-1}\sum_{\bm{\omega}\in\Omega^{l}}\norm{\bm{\varphi}_{j,\bm{\omega}}^{l}({\bm{\bm{0}}})-\overline{\bm{\varphi}}_{j}^{l}(\bm{0})}_{2}^{2}$
is the sample variance of the feature difference norm. Detailed derivation can be found in Section C of the Supplementary Material, where the Demons algorithm is established from a mathematically interpretable optimisation framework. The variance $\bm{\Sigma}_{\bm{v},j}^{l}$ of the variational posterior is predicted by a convolutional block from the concatenation of probability maps $[\widetilde{\bm{p}}_{j}^{l}\circ\phi_{j}^{<l};\widetilde{\bm{p}}^{l*}]$, where the network parameters are shared among different modalities.
@@ 4.3 Bayesian Disentangled Representation Learning
The network architecture of the decoders is given by multi-level convolutional blocks and linear upsampling. At level $l$, the output feature maps from upsampling are multiplied by the sampled categorical variables $\bm{z}^{l}$. Besides, all convolutional layers are assigned with a kernel size of 1 isotropically, to suppress spatial correlation and to form spatial equivariance. Thus, the decoded image $\widetilde{u}_{j}$ is in the same spatial orientation as the common anatomy $\bm{z}$, encouraging the registration module to identify the correct velocity fields $\bm{v}_{j}^{+}$ for reconstructing $\widehat{u}_{j}$. This connection between spatial equivariance and registration identifiability is formally established in Section E of the Supplementary Material. Furthermore, the batch normalization layers in the decoders are set as the inverse of the counterpart BN functions along the contracting path of the encoders, i.e.,
@@
$p(\bm{u}\mid\bm{z},\bm{v};\bm{\theta})$ $\propto\prod_{j=1}^{N}\prod_{\bm{\omega}\in\Omega_{j}}\exp\left\{-\frac{\absolutevalue{u_{j}(\bm{\omega})-f(\bm{z};\bm{\theta}_{j})\circ\phi_{j}^{-1}(\bm{\omega})}}{b}\right\},$ (21)
@@ 5.1 Materials
    MS-CMRSeg. This dataset provides multi-sequence cardiac MR images from 45 patients, including LGE (Late Gadolinium Enhanced), bSSFP (balanced-Steady State Free Precession), and T2-weighted scans. The images were preprocessed by affine co-registration, slice selection, resampling to $1\times 1\ \text{mm}$ and ROI extraction, giving 39, 15 and 44 slices for training, validation and test, respectively. Additional misalignment simulated by random FFDs using isotropic control point spacings of $5/10/20/40\ \text{mm}$ were made to further demonstrate registration. The ROI regions were obtained by dilating the foreground mask using a circular filter of 15-pixel radius. Evaluation was conducted on the warped manual segmentations of the myocardium, left and right ventricle. The major challenge of this dataset comes from the complex intensity patterns of the cardiac region, which can misguide the registration.
@@
[…]
@@ 5.5.2 Quantitative Results of Registration Accuracy
TABLE II: Evaluation metrics of the groupwise registration results on the MS-CMRSeg, BraTS-2021, Learn2Reg, and OASIS datasets. The top and second-best results for each dataset are highlighted in bold and underline, respectively. The ASSDs were measured in voxel units. The parameter counts are expressed in millions, and for our model there are test and training (in parentheses) values. The $p$-values were computed using the gWIs or DSCs (for the BraTS-2021 and other datasets, respectively) between the method Ours-CD and the others with a two-sided paired $t$-test. |$\absolutevalue{\det J_{\phi}\leq 0}$ | represents the proportion (in %) of voxels with negative Jacobian determinants in the predicted displacements, where the values were first calculated for the foreground region of each registered image and then averaged over all images among all test groups.
@@
None .$722\pm.104$ $3.86\pm 1.43$ N/A $<10^{-10}$ $1.430\pm 0.644$$610\pm.150$ $0.93\pm 0.42$ N/A $<10^{-10}$
[…]
CTE [5] .$766\pm.100$ $3.15\pm 1.32$ 0.154 $<10^{-10}$ $1.223\pm 0.255$$500\pm.102$ $1.27\pm 0.45$ 7.37 $<10^{-10}$
$\mathcal{X}$-CoReg [6] .$757\pm.107$ $3.31\pm 1.40$ 0.154 $<10^{-10}$ $0.728\pm 0.196$$698\pm.086$ $0.65\pm 0.17$ 7.37 $<10^{-10}$
[…]
CTE-Att .$820\pm.066$ $2.45\pm 0.74$ 8.04 $<10^{-10}$ $0.916\pm 0.210$$661\pm.094$ $0.75\pm 0.20$ 22.95 $<10^{-10}$
[…]
Ours-CN .$\bm{.842\pm.051}$ $\bm{2.17\pm 0.58}$ 5.06 (5.22) $<10^{-10}$ $\bm{0.538\pm 0.108}$$\bm{.709\pm.063}$ $\underline{0.62\pm 0.15}$ 14.91 (15.10) $<10^{-10}$
Ours-PD .$799\pm.060$ $2.74\pm 0.65$ 1.91 (2.85) $<10^{-10}$ $0.720\pm 0.151$$670\pm.072$ $0.70\pm 0.14$ 5.44 (15.06) $<10^{-10}$
Ours-CD .$\underline{.836\pm.043}$ $\underline{2.21\pm 0.47}$ 1.91 (2.85) N/A $0.663\pm 0.129$$669\pm.074$ $0.70\pm 0.15$ 5.44 (15.06) N/A
@@
DSC $\uparrow$ ASSD $\downarrow$ #Params $p$-value |$\absolutevalue{\det J_{\phi}\leq 0}$ | (%) $\downarrow$ DSC $\uparrow$ ASSD $\downarrow$ #Params $p$-value
None .$415\pm.160$ $5.27\pm 3.01$ N/A $2.4\times 10^{-4}$ N/A .$614\pm.027$ $0.96\pm 0.10$ N/A $<10^{-10}$
[…]
CTE [5] .$514\pm.373$ $6.85\pm 7.89$ 10.52 $5.1\times 10^{-2}$$2291\pm.0447$$746\pm.031$ $0.62\pm 0.09$ 8.85 $<10^{-10}$
$\mathcal{X}$-CoReg [6] .$664\pm.361$ $6.86\pm 14.1$ 10.52 $3.0\times 10^{-1}$$1330\pm.0404$$7.7$017 $0.54\pm 0.05$ 8.85 $<10^{-10}$
[…]
CTE-Att .$679\pm.069$ $2.17\pm 0.56$ 22.95 $1.4\times 10^{-2}$$0552\pm.0154$$773\pm.039$ $0.59\pm 0.13$ 22.95 $4.2\times 10^{-6}$
[…]
Ours-CD .$\underline{.793\pm.086}$ $\underline{1.39\pm 0.60}$ 5.43 (15.06) N/A .$\bm{.0066\pm.0029}$$\underline{.791\pm.008}$ $\underline{0.51\pm 0.02}$ 3.69 (13.22) N/A
Section 5.5.2 summarises the means and standard deviations of the evaluation metrics (DSC, ASSD, and gWI) on the four datasets before and after groupwise registration by different methods. The effectiveness of iterative and deep learning baselines varies across datasets. For example, all iterative approaches show inferior performance on MS-CMRSeg and Learn2Reg compared to deep learning baselines. Particularly on Learn2Reg, the evaluation reveals that APE and CTE yield marginal improvements in mean DSC, and all iterative methods even result in worse mean ASSDs after registration. In addition, deep learning baselines performed worse than APE and $\mathcal{X}$-CoReg on BraTS-2021, and all baselines (except for CTE) exhibit similar performance on OASIS. These results show the instability of both types of baselines under complex image conditions. Besides, their performance was achieved at the expense of either time-consuming iterative optimisation for each test group, or a substantial increase in parameter count for deep learning baselines.

### v3

@@ 1 Introduction
  • We propose a novel hierarchical variational auto-encoding architecture for the joint inference of latent variables. The network is able to reveal the underlying structural representations from the observations with visual semantics, based on which the intrinsic structural correspondence can be explicitly estimated in a mathematically interpretable fashion.
@@
$L$ the number of levels of latent variables NN the number of images in the group
@@ 2.1 Groupwise Image Registration
Over the past years, the shift from optimization to learning-based image registration [43, 44, 45, 46] has also motivated the development of learning-based multi-modal groupwise registration methods which predict the desired spatial transformations using neural network estimation [47, 48]. However, this deep-learning approach may inherent the same limitation of scalability from its optimisation-based counterpart. That is, the trained network for a fixed input size can hardly be utilised to register image groups of variable sizes. This hinders the potential of learning-based frameworks on large-scale groupwise registration.
@@ 2.2 Deep Feature-Based Image Registration
A closely related line of research to our work is deep feature-based image registration [49, 50, 51, 52, 53, 54]. These methods usually adopt a two-stage pipeline: in the first stage a feature extraction network is pre-trained on a surrogate task, be it segmentation [52, 53], auto-encoding [53], or contrastive learning [49, 50], followed by an instance optimisation procedure using conventional similarity measures with the learnt features in the second stage. They may have the advantage of being interpretable and capable of multi-modal registration based on modality-invariant features [49, 50, 52], compared to previous end-to-end counterparts [44, 45] that make predictions using a black-box network. However, the two-stage prediction pipeline and the requirement for ground-truth labels (segmentation annotation or aligned image pairs) limit their applicability. In contrast, Qin et al. [55] proposed using the multimodal unsupervised image-to-image translation framework [56] to reduce the multi-modal registration problem to a mono-modal one by disentangling shape and appearance representations. More recently, Deng et al. [54] proposed a unified framework where explicit modality-invariant feature extraction and unsupervised registration are learnt via an interpretable optimisation problem.
@@ 3.1 Hierarchical Bayesian Inference
$\mathcal{L}(\bm{\theta},\bm{\psi}\mid\bm{u})$ $\triangleq\begin{aligned} &\mathbb{E}_{q(\bm{z},\bm{v}\mid\bm{u};\bm{\psi})}[\log p(\bm{u}\mid\bm{z},\bm{v};\bm{\theta})]\\ &-D_{\text{KL}}\big{[}q(\bm{z},\bm{v}\mid\bm{u};\bm{\psi})\;\|\;p(\bm{z})p(\bm{v})\big{]},\end{aligned}$ (6)
@@
$D_{\text{KL}}\big{[}q(\bm{z},\bm{v}\mid\bm{u};\bm{\psi})\;\|\;p(\bm{z})p(\bm{v})\big{]}=$ $\ \mathbb{E}_{q(\bm{v}\mid\bm{u};\bm{\psi})}\left\{\sum_{l=1}^{L}\mathbb{E}_{q(\bm{z}^{<l}\mid\bm{u},\bm{v};{\bm{\psi}})}\left[D_{\text{KL}}\big{[}q(\bm{z}^{l}\mid\bm{u},\bm{v};{\bm{\psi}})\;\|\;p(\bm{z}^{l})\big{]}\right]\right\}\quad\quad\text{(i)}$ (11)
$\ +\sum_{j=1}^{N}\sum_{l=1}^{L}\mathbb{E}_{q(\bm{v}_{j}^{<l}\mid\bm{u};{\bm{\psi}})}\left[D_{\text{KL}}\big{[}q(\bm{v}_{j}^{l}\mid\bm{u},\bm{v}^{<l};{\bm{\psi}})\;\|\;p(\bm{v}_{j}^{l})\big{]}\right]\qquad\qquad\text{(ii)}$
@@ 3.2.1 Structural Representations for Variable Group Sizes
[…]
$\quad\;+(\bm{\mu}_{\bm{z},j}^{\diamond}-\bm{\mu}_{\bm{z}}^{*})^{\intercal}\bm{\Sigma}_{\bm{z},j}^{\diamond\,-1}(\bm{\mu}_{\bm{z},j}^{\diamond}-\bm{\mu}_{\bm{z}}^{*})\Bigg{]}+\text{const.}$
@@
Here, we extend the framework using categorical distribution to improve the interpretability of the structural representations. Particularly, we propose to model them by independent categorical latent variables, i.e., $q_{j}^{\diamond}(\bm{z}\mid u_{j},\bm{v}_{j};{\bm{\psi}_{j}})=\operatorname{Cat}(\bm{z};\bm{\pi}_{j,1}^{\diamond},\dots,\bm{\pi}_{j,K}^{\diamond})$, where $|\bm{\pi}_{j,k}^{\diamond}=(\pi_{\bm{\omega},j,k}^{\diamond})_{\bm{\omega}\in\Omega}\in[0,1]^{\absolutevalue{\Omega}}$ and $|\sum_{k=1}^{K}\bm{\pi}_{j,k}^{\diamond}=\bm{1}\in\mathds{R}^{\absolutevalue{\Omega}}$. Thus, the geometric mean becomes another categorical distribution
@@
$|,\bm{\pi}_{k}^{*}=\frac{\left[\prod_{j=1}^{N}\bm{\pi}_{j,k}^{\diamond}\right]^{\nicefrac{{1}}{{N}}}}{\sum_{k=1}^{K}\left[\prod_{j=1}^{N}\bm{\pi}_{j,k}^{\diamond}\right]^{\nicefrac{{1}}{{N}}}}\in[0,1]^{\absolutevalue{\Omega}}$,
@@ 3.2.2 Categorical Reparameterization Using Gumbel-Rao
Unfortunately, the Gumbel-Max trick [75] for sampling discrete random variables is not differentiable w.r.t. its parameterization, making it unsuitable for stochastic gradient estimators. Concurrent works [76, 77] have then introduced the Gumbel-Softmax (GS) distribution as a continuous relaxation of the discrete categorical variable, which admits a biased reparameterization gradient estimator. Here, we use a variant of the GS estimator, namely the Gumbel-Rao (GR) estimator [78], to further reduce the variance in gradient estimation via Rao-Blackwellisation.
@@ 4.2 Inference of Velocity Fields
[…]
is the sample variance of the feature difference norm. Detailed derivation can be found in Section C of the Supplementary Material, where the Demons algorithm is established from a mathematically interpretable optimisation framework. The variance $\bm{\Sigma}_{\bm{v},j}^{l}$ of the variational posterior is predicted by a convolutional block from the concatenation of probability maps $\widetilde{\bm{p}}_{j}^{l}\circ\phi_{j}^{<l};\widetilde{\bm{p}}^{l*}]$, where the network parameters are shared among different modalities.
@@ 4.3 Bayesian Disentangled Representation Learning
The network architecture of the decoders is given by multi-level convolutional blocks and linear upsampling. At level $l$, the output feature maps from upsampling are multiplied by the sampled categorical variables $\bm{z}^{l}$. Besides, all convolutional layers are assigned with a kernel size of 1 isotropically, to suppress spatial correlation and to form spatial equivariance. Thus, the decoded image $\widetilde{u}_{j}$ is in the same spatial orientation as the common anatomy $\bm{z}$, encouraging the registration module to identify the correct velocity fields $\bm{v}_{j}^{+}$ for reconstructing $\widehat{u}_{j}$. This connection between spatial equivariance and registration identifiability is formally established in Section E of the Supplementary Material. Furthermore, the batch normalisation layers in the decoders are set as the inverse of the counterpart BN functions along the contracting path of the encoders, i.e.,
@@
[…]
@@ 5.1 Materials
    MS-CMRSeg. This dataset provides multi-sequence cardiac MR images from 45 patients, including LGE (Late Gadolinium Enhanced), bSSFP (balanced-Steady State Free Precession), and T2-weighted scans. The images were preprocessed by affine co-registration, slice selection, resampling to $1\times 1\ \text{mm}$ and ROI extraction, giving 39, 15 and 44 slices for training, validation and test, respectively. Additional misalignment simulated by random FFDs using isotropic control point spacings of 5/10/20/40mm5/10/20/40\ $\text${mm} were made to further demonstrate registration. The ROI regions were obtained by dilating the foreground mask using a circular filter of 15-pixel radius. Evaluation was conducted on the warped manual segmentations of the myocardium, left and right ventricle. The major challenge of this dataset comes from the complex intensity patterns of the cardiac region, which can misguide the registration.
@@
    BraTS-2021. This dataset contains multi-parametric MRI scans of glioma, including native T1, T1Gd (post-contrast T1-weighted), T2-weighted, and T2-FLAIR (Fluid Attenuated Inversion Recovery). The multi-parametric MRIs of the same patients were co-registered to the same anatomical template, interpolated to the same resolution ($1\times 1\times 1\ \text{mm}$) and skull-stripped. We randomly selected 300, 50 and 150 patient cases for training, validation and test, respectively. The images were downsampled into $2\times 2\times 2\ \text{mm}$ with volume size of $80\times 96\times 80$. To demonstrate registration, four synthetic free-form deformations (FFDs) for each image with isotropic control point spacings of 5/10/15/20mm5/10/15/20\ $\text${mm} were generated to simulate misalignment. The major challenge of this task stems from the complex intensity patterns around regions of tumor, yielding ambiguous anatomical correspondences among images.
@@ 5.5.2 Quantitative Results of Registration Accuracy
TABLE II: Evaluation metrics of the groupwise registration results on the MS-CMRSeg, BraTS-2021, Learn2Reg, and OASIS datasets. The top and second-best results for each dataset are highlighted in bold and underline, respectively. The ASSDs were measured in voxel units. The parameter counts are expressed in millions, and for our model there are test and training (in parentheses) values. The $p$-values were computed using the gWIs or DSCs (for the BraTS-2021 and other datasets, respectively) between the method Ours-CD and the others with a two-sided paired $t$-test. $|\absolutevalue{\det J_{\phi}\leq 0}$ represents the proportion (in %) of voxels with negative Jacobian determinants in the predicted displacements, where the values were first calculated for the foreground region of each registered image and then averaged over all images among all test groups.
@@
None $722\pm.104$ $3.86\pm 1.43$ N/A $10^{-10}$ $1.430\pm 0.644$ $610\pm.150$ $0.93\pm 0.42$ N/A $10^{-10}$
APE [4] $746\pm.101$ $3.48\pm 1.36$ 0.154 $10^{-10}$ $0.629\pm 0.141$ $\underline{.707\pm.069}$ $\bm{0.62\pm 0.12}$ 7.37 $1.3\times 10^{-1}$
CTE [5] $766\pm.100$ $3.15\pm 1.32$ 0.154 $10^{-10}$ $1.223\pm 0.255$ $500\pm.102$ $1.27\pm 0.45$ 7.37 $10^{-10}$
$\mathcal{X}$-CoReg [6] $757\pm.107$ $3.31\pm 1.40$ 0.154 $10^{-10}$ $0.728\pm 0.196$ $698\pm.086$ $0.65\pm 0.17$ 7.37 $10^{-10}$
APE-Att $799\pm.061$ $2.78\pm 0.72$ 8.04 $10^{-10}$ $0.757\pm 0.153$ $690\pm.078$ $0.67\pm 0.14$ 22.95 $10^{-10}$
CTE-Att $820\pm.066$ $2.45\pm 0.74$ 8.04 $10^{-10}$ $0.916\pm 0.210$ $661\pm.094$ $0.75\pm 0.20$ 22.95 $10^{-10}$
Ours-PN $803\pm.062$ $2.68\pm 0.70$ 5.06 (5.22) $10^{-10}$ $\underline{0.608\pm 0.115}$ $670\pm.071$ $0.69\pm 0.13$ 14.91 (15.10) $10^{-10}$
Ours-CN $\bm{.842\pm.051}$ $\bm{2.17\pm 0.58}$ 5.06 (5.22) $10^{-10}$ $\bm{0.538\pm 0.108}$ $\bm{.709\pm.063}$ $\underline{0.62\pm 0.15}$ 14.91 (15.10) $10^{-10}$
Ours-PD $799\pm.060$ $2.74\pm 0.65$ 1.91 (2.85) $10^{-10}$ $0.720\pm 0.151$ $670\pm.072$ $0.70\pm 0.14$ 5.44 (15.06) $10^{-10}$
Ours-CD $\underline{.836\pm.043}$ $\underline{2.21\pm 0.47}$ 1.91 (2.85) N/A $0.663\pm 0.129$ $669\pm.074$ $0.70\pm 0.15$ 5.44 (15.06) N/A
@@
DSC $\uparrow$ ASSD $\downarrow$ #Params $p$-value $|\absolutevalue{\det J_{\phi}\leq 0}$ (%) $\downarrow$ DSC $\uparrow$ ASSD $\downarrow$ #Params $p$-value
None $415\pm.160$ $5.27\pm 3.01$ N/A $2.4\times 10^{-4}$ N/A $614\pm.027$ $0.96\pm 0.10$ N/A $10^{-10}$
APE [4] $554\pm.384$ $6.14\pm 7.66$ 10.52 $9.7\times 10^{-2}$ $5297\pm.0814 .$$777\pm.030$ $0.59\pm 0.09$ 8.85 $7.9\times 10^{-8}$
CTE [5] $514\pm.373$ $6.85\pm 7.89$ 10.52 $5.1\times 10^{-2}$ $2291\pm.0447 .$$746\pm.031$ $0.62\pm 0.09$ 8.85 $10^{-10}$
$\mathcal{X}$-CoReg [6] $664\pm.361$ $6.86\pm 14.1$ 10.52 $3.0\times 10^{-1}$ $1330\pm.0404 .$$777\pm.017$ $0.54\pm 0.05$ 8.85 $10^{-10}$
APE-Att $687\pm.092$ $09\pm 0.57 22.$95 $2.9\times 10^{-3}$ $0479\pm.0130 .$$777\pm.018$ $0.57\pm 0.05$ 22.95 $10^{-10}$
CTE-Att $679\pm.069$ $17\pm 0.56 22.$95 $1.4\times 10^{-2}$ $0552\pm.0154 .$$773\pm.039$ $0.59\pm 0.13$ 22.95 $4.2\times 10^{-6}$
Ours-CN $\bm{.803\pm.084}$ $\bm{1.32\pm 0.57}$ 14.90 (15.10) $9.3\times 10^{-2}$ $1746\pm.0305 .$$\bm{.803\pm.007}$ $\bm{0.49\pm 0.02}$ 13.16 (13.26) $10^{-10}$
Ours-CD $\underline{.793\pm.086}$ $\underline{1.39\pm 0.60}$ 5.43 (15.06) N/A $\bm{.0066\pm.0029}$ $\underline{.791\pm.008}$ $\underline{0.51\pm 0.02}$ 3.69 (13.22) N/A
Table II summarises the means and standard deviations of the evaluation metrics (DSC, ASSD, and gWI) on the four datasets before and after groupwise registration by different methods. The effectiveness of iterative and deep learning baselines varies across datasets. For example, all iterative approaches show inferior performance on MS-CMRSeg and Learn2Reg compared to deep learning baselines. Particularly on Learn2Reg, the evaluation reveals that APE and CTE yield marginal improvements in mean DSC, and all iterative methods even result in worse mean ASSDs after registration. In addition, deep learning baselines performed worse than APE and $\mathcal{X}$-CoReg on BraTS-2021, and all baselines (except for CTE) exhibit similar performance on OASIS. These results show the instability of both types of baselines under complex image conditions. Besides, their performance was achieved at the expense of either time-consuming iterative optimisation for each test group, or a substantial increase in parameter count for deep learning baselines.